OpenStudy (anonymous):

what is the square root of 56

OpenStudy (arabpride):

\(\huge\color{green}{Hey~there!}\) \(\huge\color{blue}{Please~ask~in~math~section!}\)

OpenStudy (anonymous):

solve this question 3x+5x+2=28

OpenStudy (anonymous):

Simplifying 3x + 5x + 2 = 28 Reorder the terms: 2 + 3x + 5x = 28 Combine like terms: 3x + 5x = 8x 2 + 8x = 28 Solving 2 + 8x = 28 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + 8x = 28 + -2 Combine like terms: 2 + -2 = 0 0 + 8x = 28 + -2 8x = 28 + -2 Combine like terms: 28 + -2 = 26 8x = 26 Divide each side by '8'. x = 3.25 Simplifying x = 3.25 See similar equations: Equations solver categories Equations solver - equations involving one unknown Quadratic equations solver Percentage Calculator - Step by step Derivative calculator - step by step Greatest Common Factor Least Common Multiple System of equations - step by step solver Simplifying 3x + 5x + 2 = 28 Reorder the terms: 2 + 3x + 5x = 28 Combine like terms: 3x + 5x = 8x 2 + 8x = 28 Solving 2 + 8x = 28 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + 8x = 28 + -2 Combine like terms: 2 + -2 = 0 0 + 8x = 28 + -2 8x = 28 + -2 Combine like terms: 28 + -2 = 26 8x = 26 Divide each side by '8'. x = 3.25 Simplifying x = 3.25 See similar equations: Equations solver categories Equations solver - equations involving one unknown Quadratic equations solver Percentage Calculator - Step by step Derivative calculator - step by step Greatest Common Factor Least Common Multiple System of equations - step by step solver

OpenStudy (anonymous):

56 = 2^3 * 7 √56 = √(4*14) = √4*√14 = 2√14 This is as simplified as it can be while still being an exact answer. It is approximately 7.483 in decimal form.

OpenStudy (k_lynn):

You might want to try posting math questions in the maths section. And square root means what number times itself is equal to whatever number. In your case, it will look like this \[\sqrt{56}\] So what number times itself equals 56. It won't be a whole number. If you have a scientific calculator, you can find out the answer. There's also one on the computer if you need it.